# Dynamics of small bubbles in turbulence in non-dilute conditions

### v1

@@ I Introduction
This work focuses on the dynamics of bubbles that are small with respect to the Kolmogorov scale of the turbulent flow (also referred to as microbubbles), are undeformable and are subject to strong turbulence so that gravity plays a negligible role. Semantically, this coincides with the high Froude number, low Weber number limit. In numerical simulations, such bubbles have traditionally been studied in a one-way coupling paradigm, meaning that only the advection from the fluid phase exerted onto the bubble phase is taken into account. Intuitively, such an approximation is justified if the concentration of bubbles is dilute, such that both the collective effect of the back-reaction of the bubbles onto the fluid as well as interactions between bubbles themselves can be neglected. Once the bubble volume fraction increases, however, this approximation is challenged, and more sophisticated methods that take all such interactions into account are warranted, potentially exacerbating the numerical costs. Recently, we have introduced an efficient method for simulating small particles in turbulence in the so-called four-way coupling regime, taking the two-sided momentum exchange between particle and fluid phase into account, as well as excluded-volume interactions between the particles, at only a very limited additional numerical cost with respect to one-way coupling [11].
@@ II Numerical set-up
The flow field $\bm{u}(\bm{r},t)$ in the fluid phase in Eulerian space $\bm{r}$ and time $t$ is governed by the conservation of momentum and mass as
$\frac{\mathrm{D}\bm{u}}{\mathrm{D}t}=-\bm{\nabla}p+\nu\Delta\bm{u}+\bm{f}+\bm{f}_{p},$ (1a)
$\bm{\nabla}\cdot\bm{u}=0,$ (1b)
where $p$ denotes the pressure and $\nu$ is the kinematic viscosity. The flow is driven by an external force $\bm{f}$, acting on forcing wavenumber $k_{f}$, defined in Fourier space as $\hat{\bm{f}}(\bm{k})=\varepsilon\hat{\bm{u}}(\bm{k})/\sum_{k_{f}\leq|\bm{k}|<k_{f}+1}|\hat{\bm{u}}(\bm{k})|^{2}$, ensuring smoothness and a constant energy injection rate $\varepsilon$ [12].
@@
Finally, the particle phase is evolved through Lagrangian interpolation following the Maxey-Riley equations [15], where we retain only the dominant forces, being the pressure gradient force, added mass and Stokes drag, yielding
$\frac{\mathrm{d}\bm{v}}{\mathrm{d}t}$ $=\beta\frac{\mathrm{D}\bm{u}(\bm{x},t)}{\mathrm{D}t}-\frac{1}{\tau_{p}}(\bm{v}-\bm{u}(\bm{x},t)),$ (3a)
$\frac{\mathrm{d}\bm{x}}{\mathrm{d}t}$ $=\bm{v}+\textrm{(collisions)}.$ (3b)
Here, $\beta\equiv 3/(1+2\rho_{p}/\rho_{f})$ and $\tau_{p}$ is the reciprocal Stokes drag coefficient or particle response time, which is typically compared with the Kolmogorov time $\tau_{\eta}\equiv\sqrt{\nu/\varepsilon}$ of the turbulence to yield the Stokes number $\textrm{St}\equiv\tau_{p}/\tau_{\eta}$. Finally, we resolve collisions between particles by enforcing the excluded-volume interactions, assuming spherical particles with collision diameter $D$, using the ‘You Only Collide Once’ (YOCO) algorithm, see Ref. [11].
This method presumes that collisions happen instantaneously. In reality, there will be an elastic time scale involved in the bubble collision process. Our method thus represents the limit in which this elastic time scale of collisions is much smaller than the smallest dynamical time scale of the turbulent flow, represented by $\tau_{\eta}$.
@@ III Bubble-fluid interaction
To understand the effect of the four-way coupled bubbles on the underlying turbulent flow, we start by studying how they affect the kinetic energy. Changes in the energetic properties of the flow will be a consequence of the back-reaction that the bubbles exert on the fluid phase. In Fig. 2, we show the total kinetic energy $E_{\textrm{tot}}$ (panel a) in the fluid phase for the four-way coupled system at varying bubble volume fraction, compared with its one-way coupled counterpart, as well as the corresponding spectral distribution of energy $E(k)$ as a function of wavenumber $k$ (panel b).
@@
It shows that only when one goes to extremely high volume fraction of $\alpha\gtrsim 10\%$, an appreciable effect on the total kinetic energy is obtained. However, even at these very high volume fractions, the quantitative effect on total kinetic energy remains very small, at the sub-5% level. We furthermore point out that the effect on the kinetic energy collapses for different $D$, which is a consequence of the proportionality of the feedback force of each bubble with the bubble volume $\mathcal{V}_{p}$, making the total feedback force effectively proportional to the local volume fraction, see Eq. (2). Considering the different spectral scales, we also find just small effects, which only become clear at the highest volume fractions. We notice an increase of energy at the very smallest scales (highest $k$), corresponding to the scale of the bubbles, followed by a slight attenuation of spectral energy in the dissipative range and an even slighter increase of energy in the forcing and inertial scales. This is in agreement with earlier results for two-way coupled bubbles [19], and can be understood from basic arguments: if one approximates the feedback force as $\bm{f}_{p}\approx\alpha\frac{\mathrm{D}\bm{u}}{\mathrm{D}t}$, see Eq. (2), then a substitution into the Navier-Stokes Eq. (1a), reveals that both the energy injection rate $\varepsilon$ as well as the viscosity $\nu$ are effectively enhanced by a factor $1/(1-\alpha)$, which slightly increases the energy in the inertial range and moves the Kolmogorov scale to lower $k$, thereby decreasing the energetic content in the dissipative range.
The relatively small magnitude of the effect of the four-way coupled bubbles on the fluid kinetic energy can be understood from the nature of the back-reaction force itself. For fluids laden with heavy particles, strong turbulence attenuation is obtained even at only moderate volume fractions [20, 21, 22, 23], see also the inset in panel (a). The massless bubbles, on the contrary, do not carry any momentum themselves. As a consequence, the only back-reaction in the case of bubbles is due to their added mass, i.e. the first term in the square brackets in Eq. (2), while the second term vanishes, yielding only a relatively small force on the underlying flow.
@@ IV Preferential sampling and clustering
The results are shown in Fig. 3 for varying volume fraction $\alpha$ and for different bubble sizes $D$ and turbulence Taylor-scale Reynolds number $\textrm{Re}_{\lambda}$. The figure shows a clear transition between a dilute regime at low $\alpha$ and a denser regime at higher $\alpha$. In the dilute regime, the preferential sampling of the four-way coupled bubbles is indistinguishable from that of the one-way coupled bubbles. In the denser regime, preferential sampling decreases as bubbles start to interact through their excluded-volume and eventually approaches the Eulerian average expected for uniformly distributed bubbles at the highest volume fractions. The transition between the two regimes occurs at around a critical volume fraction $\alpha=\alpha_{c}\approx 1\%$, independent of the bubble size $D$ in the considered range, as indicated by the collapse in panel (a). We recall that this collapse is with respect to the volume fraction $\alpha$, which for each bubble size corresponds to a different number of bubbles. The small excess oversampling of enstrophy for the largest bubbles ($D=1.63\eta$) at low $\alpha$ beyond the one-way coupled bubbles is attributed to the self-induced vorticity sensed by the bubbles due to their back-reaction on the fluid.
@@
To assess the Re-dependence of this transition point, we consider the preferential sampling for different $\textrm{Re}_{\lambda}$ in panel (b). This yields different levels of preferential sampling due to their different levels of intermittency, but we can linearly rescale the enstrophy to be between 0 and 1 (inset). In panel (c) we show that these curves then collapse well upon rescaling the volume fraction by a factor $\textrm{Re}_{\lambda}^{-1/2}$. This suggests that the transition point, associated with the volume fraction of vortex filaments, scales as $\alpha_{c}\propto\textrm{Re}_{\lambda}^{-1/2}$. However, it must be noted that the accessible dynamic range of $\textrm{Re}_{\lambda}$ is limited.
@@
As we approach this transition point at which the most preferential regions become filled, we should also expect the morphology of the bubble clusters to change. To investigate this, we compute the fractal covering dimension $d$ of the total set of bubbles, capturing the spatial dimensionality of the bubble configuration. Details on this computation are provided in Appendix A. The resulting fractal dimensions are shown in Fig. 4. This indeed shows how the bubble cluster conformation changes from one dimensional ($d\approx 1$) line-like structures at low $\alpha$, following the morphology of vortex filaments, to a gradually more volumetric conformation at higher $\alpha$, approaching full three dimensionality ($d\to 3$). This change can also be appreciated visually from the example snapshot in Fig. 1. For the fractal covering dimension, there is an explicit dependence on $D$, as larger bubbles remain in a line-like configuration till higher $\alpha$ than smaller bubbles when filling up the vortex filaments. This is also evident from the apparent collapse when rescaling the abscissa by the bubble size $D$ (inset).
@@ V Lagrangian velocity structure function
The local scaling exponents $\zeta_{p}(\tau)$ of the Lagrangian velocity structure functions are known to show a notable decrease around the dissipative scale $\tau=\mathcal{O}(\tau_{\eta})$, before plateauing into the inertial range at $\tau>\mathcal{O}(10\tau_{\eta})$ [38]. This so-called ‘viscous dip’ is associated with trapping in vortex filaments [39, 40] and is hence amplified for light particles [41].
In Fig. 5, we show the obtained ESS Lagrangian velocity structure function scaling exponents of the 4th order for the four-way coupled bubbles, alongside the one-way coupled bubbles and tracer particles for comparison. It shows that at low volume fraction $\alpha\lesssim 1\%$, the scaling exponents closely follow that of the one-way coupled bubbles, while at larger volume fractions $\alpha\gtrsim 1\%$ the scaling exponents get closer to that of tracers. At these higher volume fractions, preferential sampling and thus trapping in vortex filaments decreases, resulting in a less pronounced viscous dip. At larger time scales, approaching the inertial range $\tau\gtrsim 10\tau_{\eta}$, all scaling exponents collapse within error bars. On the other hand, at the shortest time scales $\tau\ll\tau_{\eta}$, while one-way coupled bubbles and tracers recover the smooth solution $\lim_{\tau\to 0}\zeta_{4}(\tau)\approx 2$, the four-way coupled bubbles approach the rough, shock-like solution $\lim_{\tau\to 0}\zeta_{4}(\tau)\approx 1$ due to the effect of the instantaneous collisions, akin to Burgers turbulence [42, 43]. These collisions dominate the statistics of velocity differences at the shortest time scales even for the smallest volume fractions when collisions are increasingly rare. However, crucially, as can be seen from the figure, the effect of collisions quickly vanishes as $\tau$ approaches $\tau_{\eta}$ and beyond. This indicates that results obtained for one-way coupled bubbles are correctly approaching those for four-way coupled bubbles in the dilute limit $\alpha\lesssim 1\%$ at all relevant dynamical time scales when it comes to the Lagrangian velocity structure function. In the non-dilute regime $\alpha\gtrsim 1\%$, the scaling exponents are well-described by a simple linear combination of the scaling exponents of tracers and one-way coupled bubbles $\zeta_{4}^{\textrm{(fit)}}=\gamma\zeta_{4}^{\textrm{(tracer)}}+(1-\gamma)\zeta_{4}^{\textrm{(one-way)}}$, with fit parameter $\gamma$, as indicated by the green dashed lines in Fig. 5.
@@ VI Pair dispersion
$\Delta x_{ij}(\tau)\equiv|\bm{x}_{i}(t+\tau)-\bm{x}_{j}(t+\tau)|,$ (6)
which initially are close together $0.9D<\Delta x_{ij}(0)<3.5\eta$. In Fig. 6, we plot the mean squared displacement $\langle\Delta x_{ij}^{2}(\tau)\rangle$ for four-way coupled bubbles at different volume fractions as well as for tracers and one-way coupled bubbles for comparison. The pair dispersion shows an initial exponential growth, followed by the famous Richardson superdiffusive growth as $\propto\tau^{3}$ in the inertial range, and a diffusive regime $\propto\tau$ at long time beyond the integral scale ($\approx 10^{2}\tau_{\eta}$) [44, 45, 46]. Note that the tracers, one-way, and four-way coupled bubbles exhibit similar mean dispersion, and there are no measurable differences between the one-way and four-way coupled bubbles within the statistical accuracy. This indicates that the collisions between bubbles do not alter their mean dispersive behavior.
@@ VII Conclusions and outlook
Our findings show that the question whether or not four-way coupling is important in bubbly turbulence is a subtle question to answer and is intrinsically dependent on the type of observable that one is interested in. While we have shown that some observables are virtually insensitive to four-way coupling at all volume fractions, other observables show a dilute, unaffected regime and a denser, affected regime, while still other observables show effects of four-way coupling at all considered volume fractions.

### v2

@@ I Introduction
This work focuses on the dynamics of bubbles that are small with respect to the Kolmogorov scale of the turbulent flow (also referred to as microbubbles), are undeformable and are subject to strong turbulence so that gravity plays a negligible role. Semantically, this coincides with the high Froude number, low Weber number limit. In numerical simulations, such bubbles have traditionally been studied in a one-way coupling paradigm, meaning that only the advection from the fluid phase exerted onto the bubble phase is taken into account. Such one-way coupling simulations are able to reproduce phenomena such as preferential concentration, dispersion and clustering [39, 21, 4, 43, 18, 54, 5]. Intuitively, this approximation is justified if the concentration of bubbles is dilute, such that both the collective effect of the back-reaction of the bubbles onto the fluid as well as interactions between bubbles themselves can be neglected. Once the bubble volume fraction increases, however, this approximation is challenged, and more sophisticated methods that take all such interactions into account are warranted. This can be achieved using methods that fully resolve the geometry of both phases and their boundary conditions, such as the Immersed Boundary Method [48, 27, 46, 52, 56], or using approximate point-based methods such as Refs. [32, 57]. While fully resolving methods offer increased accuracy, they exacerbate the computational cost, making it intractable to scale up to simulations of millions of small particles or bubbles. Recently, we have introduced an efficient novel point-based method for simulating small particles in turbulence in the so-called four-way coupling regime, taking the two-sided momentum exchange between particle and fluid phase into account, as well as excluded-volume interactions between the particles, at only a very limited additional numerical cost with respect to one-way coupling [23].
@@ II Numerical set-up
The flow field $\bm{u}(\bm{r},t)$ in the fluid phase in Eulerian space $\bm{r}$ and time $t$ is governed by the conservation of momentum and mass as {subequations} {gather} D uD t = - p + νΔu + f + f_p,
u=0, where $p$ denotes the pressure and $\nu$ is the kinematic viscosity. The flow is driven by an external force $\bm{f}$, acting on forcing wavenumber $k_{f}$, defined in Fourier space as $\hat{\bm{f}}(\bm{k})=\varepsilon\hat{\bm{u}}(\bm{k})/\sum_{k_{f}\leq|\bm{k}|<k_{f}+1}|\hat{\bm{u}}(\bm{k})|^{2}$, ensuring smoothness and a constant energy injection rate $\varepsilon$ [36].
@@
Finally, the particle phase is evolved through Lagrangian interpolation following the Maxey-Riley equations [40], where we retain only the dominant forces, being the pressure gradient force, added mass and Stokes drag, yielding {subequations} {align} d vd t = βD u(x,t)D t -1τp(v-u(x,t)),
d xd t = v + (collisions). Here, $\beta\equiv 3/(1+2\rho_{p}/\rho_{f})$ and $\tau_{p}$ is the reciprocal Stokes drag coefficient or particle response time, which is typically compared with the Kolmogorov time $\tau_{\eta}\equiv\sqrt{\nu/\varepsilon}$ of the turbulence to yield the Stokes number $\textrm{St}\equiv\tau_{p}/\tau_{\eta}$.
We resolve collisions between particles by assuming hard spherical particles with collision diameter $D$, using the ‘You Only Collide Once’ (YOCO) algorithm. This method performs elastic collisions on particles that are forecast to collide in the next time step and re-enforces excluded-volume for particles that are overlapping, treating the most severe overlap or collision for each particle at every time step, see Ref. [23] for a detailed description. This method presumes that collisions happen instantaneously. In reality, there will be an elastic time scale involved in the bubble collision process. Our method thus represents the limit in which this elastic time scale of collisions is much smaller than the smallest dynamical time scale of the turbulent flow, represented by $\tau_{\eta}$.
@@ III Bubble-fluid interaction
To understand the effect of the four-way coupled bubbles on the underlying turbulent flow, we start by studying how they affect the kinetic energy. Changes in the energetic properties of the flow will be a consequence of the back-reaction that the bubbles exert on the fluid phase. In Fig. 2, we show the total kinetic energy $E_{\textrm{tot}}$ (panel a) in the fluid phase for the four-way coupled system at varying bubble volume fraction, compared with its one-way coupled counterpart (with the same $\beta=3$ and $\textrm{St}=1$), as well as the corresponding spectral distribution of energy $E(k)$ as a function of wavenumber $k$ (panel b).
@@
[…]
@@ IV Preferential sampling and clustering
The results are shown in Fig. 3 for varying volume fraction $\alpha$ and for different bubble sizes $D$ and turbulence Taylor-scale Reynolds number $\textrm{Re}_{\lambda}$. The figure shows a clear transition between a dilute regime at low $\alpha$ and a denser regime at higher $\alpha$. In the dilute regime, the preferential sampling of the four-way coupled bubbles is indistinguishable from that of the one-way coupled bubbles. In the denser regime, preferential sampling decreases as bubbles start to interact through their excluded-volume and eventually approaches the Eulerian average expected for uniformly distributed bubbles at the highest volume fractions. The transition between the two regimes occurs at around a critical volume fraction $\alpha=\alpha_{c}\approx 1\%$, independent of the bubble size $D$ in the considered range, as indicated by the collapse in panel (a). We recall that this collapse is with respect to the volume fraction $\alpha$, which for each bubble size corresponds to a different number of bubbles. The largest bubbles ($D=1.63\eta$) show a small excess oversampling of enstrophy beyond the one-way coupled bubbles at low $\alpha$. This is attributed to the self-induced vorticity sensed by the bubbles that results from the perturbations produced by the localized back-reaction $\bm{f}_{p}$ on the fluid.
@@
To assess the Re-dependence of this transition point, we consider the preferential sampling for different $\textrm{Re}_{\lambda}$ in panel (b). This yields different levels of preferential sampling due to their different levels of intermittency, but we can linearly rescale the enstrophy to be between 0 and 1 (inset). In panel (c) we show that these curves then collapse well upon rescaling the volume fraction by a factor $\textrm{Re}_{\lambda}^{-1/2}$. This suggests that the transition point, associated with the volume fraction of vortex filaments, scales as $\alpha_{c}\propto\textrm{Re}_{\lambda}^{-1/2}$, although it must be noted that the accessible dynamic range of $\textrm{Re}_{\lambda}$ is limited. We point out that this scaling relation is an empirical finding here, and how to rationalize it through scaling arguments remains an interesting open question. This would relate to both the size as well as the effective number density of vortex filaments, and would possibly also be affected by the lifetime of vortex filaments. The approach demonstrated in this work using four-way coupled bubbles seems to provide a promising pathway to study such questions.
@@
As we approach this transition point at which the most preferential regions become filled, we should also expect the morphology of the bubble clusters to change. To investigate this, we compute the fractal covering dimension $d$ of the total set of bubbles, capturing the spatial dimensionality of the bubble configuration. Details on this computation are provided in Appendix A. The resulting fractal dimensions are shown in Fig. 4. This indeed shows how the bubble cluster conformation changes from one dimensional ($d\approx 1$) line-like structures at low $\alpha$, following the morphology of vortex filaments, to a gradually more volumetric conformation at higher $\alpha$, approaching full three dimensionality ($d\to 3$). This change can also be appreciated visually from the example snapshot in Fig. 1. For the fractal covering dimension, there is an explicit dependence on $D$, as larger bubbles remain in a line-like configuration until higher $\alpha$ than smaller bubbles when filling up the vortex filaments. This is also evident from the apparent empirical collapse when rescaling the abscissa by the bubble size $D$ (inset). The fractal dimension going slightly below the one-way coupling line for low $\alpha$ is attributed to the finite number of bubbles, leading to point-like contributions from individual bubbles at the lowest $\alpha$, as motivated in Appendix A.
@@ V Lagrangian velocity structure function
The resulting local scaling exponents $\zeta_{p}(\tau)$ of the Lagrangian velocity structure functions are known to show a notable decrease around the dissipative scale $\tau=\mathcal{O}(\tau_{\eta})$, before plateauing into the inertial range at $\tau>\mathcal{O}(10\tau_{\eta})$ [10]. This so-called ‘viscous dip’ is associated with trapping in vortex filaments [12, 6] and is hence amplified for light particles [42].
In Fig. 5, we show the obtained ESS Lagrangian velocity structure function scaling exponents of the 4th order for the four-way coupled bubbles, alongside the one-way coupled bubbles and tracer particles for comparison. It shows that at low volume fraction $\alpha\lesssim 1\%$, the scaling exponents closely follow that of the one-way coupled bubbles, while at larger volume fractions $\alpha\gtrsim 1\%$ the scaling exponents get closer to that of tracers. At these higher volume fractions, preferential sampling and thus trapping in vortex filaments decreases, resulting in a less pronounced viscous dip. At larger time scales, approaching the inertial range $\tau\gtrsim 10\tau_{\eta}$, all scaling exponents collapse within error bars. On the other hand, at the shortest time scales $\tau\ll\tau_{\eta}$, while one-way coupled bubbles and tracers recover the smooth solution $\lim_{\tau\to 0}\zeta_{4}(\tau)\approx 2$, the four-way coupled bubbles approach the rough, shock-like solution $\lim_{\tau\to 0}\zeta_{4}(\tau)\approx 1$ due to the effect of the instantaneous collisions, akin to Burgers turbulence [16, 22]. Since in this short time limit, no other dynamical processes are active, these collisions quickly dominate the statistics of velocity differences at the shortest time scales even for the smallest volume fractions when collisions are increasingly rare. However, crucially, as can be seen from the figure, the effect of collisions quickly vanishes as $\tau$ approaches $\tau_{\eta}$ and beyond. This indicates that results obtained for one-way coupled bubbles are correctly approaching those for four-way coupled bubbles in the dilute limit $\alpha\lesssim 1\%$ at all relevant dynamical time scales when it comes to the Lagrangian velocity structure function. In the non-dilute regime $\alpha\gtrsim 1\%$, the scaling exponents are well-described by a simple linear combination of the scaling exponents of tracers and one-way coupled bubbles $\zeta_{4}^{\textrm{(fit)}}=\gamma\zeta_{4}^{\textrm{(tracer)}}+(1-\gamma)\zeta_{4}^{\textrm{(one-way)}}$, with fit parameter $\gamma$, as indicated by the green dashed lines in Fig. 5.
@@ VI Pair dispersion
$x_{ij}(\tau)\equiv|\bm{x}_{i}(t+\tau)-\bm{x}_{j}(t+\tau)|,$ (4)
which initially are close together $0.9D<x_{ij}(0)<3.5\eta$. To eliminate the effect of the initial condition, we study the quantity
$\Delta x_{ij}(\tau)\equiv x_{ij}(\tau)-x_{ij}(0).$ (5)
In Fig. 6, we plot the mean squared displacement $\langle\Delta x_{ij}^{2}(\tau)\rangle$ and its scaling behavior for four-way coupled bubbles at different volume fractions as well as for tracers and one-way coupled bubbles for comparison. The pair dispersion shows an initial ballistic growth, followed by the famous Richardson superdiffusive growth as $\propto\tau^{3}$ in the inertial range, and a diffusive regime $\propto\tau$ at long time beyond the integral scale ($\approx 10^{2}\tau_{\eta}$) [50, 51, 5]. Note that the tracers, one-way, and four-way coupled bubbles exhibit similar mean dispersion. Apart from small differences at short time scales $\sim\mathcal{O}(\tau_{\eta})$ the one-way and four-way coupled bubbles largely collapse within the statistical accuracy. This indicates that the collisions between bubbles do not strongly alter their mean dispersive behavior.
@@ VII Conclusions and outlook
Our findings show that the question whether or not four-way coupling is important in bubbly turbulence is a subtle question to answer and is intrinsically dependent on the type of observable that one is interested in. While we have shown that some observables are virtually insensitive to four-way coupling at all volume fractions, other observables show a dilute, unaffected regime and a denser, affected regime, while still other observables show effects of four-way coupling at all considered volume fractions. Note, however, that we have here considered massless bubbles with $\textrm{St}=1$, which show the strongest effect of preferential concentration. Moving to different St, higher or lower, will naturally result in a diminishing effect of preferential concentration and hence also the importance of the effects of four-way coupling observed here are expected to decrease as local densities decrease. A rigorous study of how preferential concentration depends on St and density contrast $\beta$ is provided in Ref. [18].